OpenStudy (anonymous):

the observation deck of a tall skyscraper 370 m above the street. Determine the time required for a penny to free fall

OpenStudy (anonymous):

X = Xo + Vo*t + 0.5*g*t^2 Let's take Xo to be 0. This means X is 370. g = 9.8 m/s^2 The penny is at rest to begin with, so Vo = 0. 370 = 0 + 0 + 0.5*9.8*t^2 370 = 4.9*t^2 75.51 = t^2 t = 8.7 seconds